### v1

@@ 2.2 Detection with LISA
While traditional radio pulsar searches are limited in detecting short-period binaries due to orbital acceleration effects (see bottom right panel of Fig. 1), LISA will offer a complementary window into the MSP binary population. In particular, LISA will be sensitive to systems with orbital periods below one hour across the entire Milky Way (Amaro-Seoane et al., 2023). In the case of circular binaries, as considered here due to the orbit being circularised during the recycling phase via stable mass transfer, the system emits continuous, quasi-monochromatic gravitational waves. The gravitational-wave signal from such a binary can be described by eight parameters:
@@
where $\mathcal{A}$ is the gravitational wave amplitude, $f_{\rm GW}=2/P_{\rm orb}$ is the gravitational wave frequency (twice the orbital frequency), and $\dot{f}_{\rm GW}$ is its time derivative. The parameters $(\lambda,\beta)$ denote the ecliptic longitude and latitude, respectively; $\iota$ is the inclination angle of the orbital plane; $\psi$ is the polarisation angle; and $\phi_{0}$ is the initial phase of the waveform (e.g. LISA Consortium Waveform Working Group et al., 2023). The gravitational wave amplitude is given by
@@
where $G$ is the gravitational constant, $c$ is the speed of light, and $d$ is the distance to the binary. This shows that the gravitational wave amplitude depends on the binary’s frequency, distance, and chirp mass $\mathcal{M}$, defined as
@@
where $T_{\rm obs}$ is the observation time and $S_{n}(f)$ is the LISA noise power spectral density. In the following analysis, we assume the LISA noise power spectral density as given in the LISA Science Requirements Document (SciRD, Babak et al., 2021). To assess a binary’s detectability, one can invert Eq. (5), setting the observation time and a detection threshold.
Figure 2 illustrates the detectable region in the companion mass–gravitational-wave frequency parameter space, assuming an MSP mass of $1.4\leavevmode\nobreak\ \mathrm{M}_{\odot}$ and distances of 5 (dark grey), 8 (grey), and 11 kpc (light grey). These distances span the inner Milky Way, which is relevant for the observed $\gamma$-ray excess (e.g. Hooper & Slatyer, 2013). The figure shows that LISA detectability at these distances requires frequencies of at least $\sim$1 mHz (corresponding to orbital periods of less than $\sim$30 min). For lower companion masses, the required frequency increases: approximately 3 mHz for $\sim$0.1 M, and up to $\sim$$10^{-2}$ Hz for $\sim$0.02 M.
Figure 2: Detectability with LISA of a binary system containing a 1.4 M MSP, shown as a function of companion mass and gravitational-wave frequency (bottom x-axis) or orbital period (top x-axis). Grey contours indicate LISA’s 4-year signal-to-noise threshold ($\rho_{\rm 4\,yr}=7$) for sources located at 5 (dark grey), 8 (grey), and 11 kpc (light grey). Dashed horizontal lines mark approximate boundaries between different classes of MSP binaries: black widows (low-mass companions, $M\lesssim 0.06$ M), redbacks and He WDs (intermediate-mass companions, $0.06\lesssim M\lesssim 0.5$ M), and more massive donors such as CO/ONeMg WDs or He stars ($M\gtrsim 0.5$ M). The grey solid curve marks the Roche-lobe overflow (RLOF) boundary for WD donors, indicating the onset of stable mass transfer. Observed MSP binaries are shown as coloured symbols, following the same classification as in Fig. 1. UCXBs lying near the WD RLOF line are shown as light blue pentagons. The two-segment grey arrow illustrates the expected evolutionary track of close binaries: initial orbital decay due to gravitational-wave emission drives the system toward higher frequencies (rightward), until reaching a period minimum at RLOF; mass transfer then leads to lower companion masses and longer periods (downward and leftward). Detached binaries are therefore not expected to exist beyond the RLOF line.
While the detection of NS binaries in the inner 3 kpc is, in principle, possible at gravitational-wave frequencies above a few millihertz, the companion star may—depending on its mass—fill its Roche lobe and initiate mass transfer before the system reaches the higher-frequency part of the LISA sensitivity window. As a reference, we plot in Fig. 2 the Roche lobe–filling radius of a WD donor using the commonly adopted zero-temperature radius–mass relation from Verbunt & Rappaport (1988):
@@
Binary evolution simulations show that MSPs with He WD companions initiate mass transfer at orbital periods of 5–7 minutes (e.g., Sengar et al., 2017), after which the orbit begins to widen as the donor expands in response to mass loss. This sets an upper frequency limit for detecting such systems in the detached phase. In contrast, simulations by Bobrick et al. (2017) show that systems with CO WD or ONeMg WD companions undergo unstable mass transfer when reaching these short orbital periods, leading to the tidal disruption of the donor. Systems with He-star companions—i.e. stripped stellar cores still burning helium and thus larger and less compact than degenerate WDs—with masses up to approximately $1.2\,\mathrm{M}_{\odot}$ can initiate stable mass transfer at orbital periods between 10 and 80 minutes, essentially set by the Roche lobe radius of the He star (Götberg et al., 2020; Wang et al., 2021). The exact minimum period depends on the He star’s mass, with shorter periods corresponding to physically smaller (lower-mass) donors. Like systems with He WD donors, these binaries evolve toward longer orbital periods after reaching the minimum. This characteristic evolution—initial inspiral driven by gravitational-wave emission, followed by mass transfer and orbital widening—is schematically illustrated with a grey two-segment arrow in Fig. 2. More massive He stars (above $1.2\,\mathrm{M}_{\odot}$) tend to initiate mass transfer that is too rapid to remain stable, often resulting in unstable mass transfer or mergers.
@@ 3 Population synthesis of MSP binaries in the Milky Way bulge
Many theories have been proposed to explain the origin of the Milky Way’s bulge (e.g. Nataf, 2017; Shen & Li, 2016), each with distinct implications for the underlying stellar populations and binary formation channels. Here, we consider two contrasting scenarios: one in which the bulge formed in situ as part of the Milky Way disc through internal evolutionary processes, and another in which it assembled through the accretion and disruption of globular clusters. While these scenarios may be somewhat simplified, they provide a useful framework for exploring the potential properties and detectability of MSP binary populations in the bulge.
In the accretion scenario, the main idea is that, due to dynamical friction, globular clusters gradually spiral toward the Galactic Centre. In the inner few kiloparsecs, the tidal field of the Galaxy can exceed a cluster’s self-gravity, stripping stars and potentially dissolving the cluster entirely (e.g. Tremaine et al., 1975; Capuzzo-Dolcetta, 1993; Antonini et al., 2012; Arca-Sedda & Capuzzo-Dolcetta, 2014; Gnedin et al., 2014). The globular clusters observed today are therefore likely the survivors of this process, representing only a small fraction of the original population. Globular clusters in the Milky Way halo are also known to be overabundant in MSPs (see Section 2). It is therefore reasonable to hypothesise that the MSPs responsible for the $\gamma$-ray excess could have originated in dense star clusters in the halo, which were subsequently disrupted, depositing their MSPs into the Galactic bulge. Brandt & Kocsis (2015) showed that, in this scenario, the amplitude, angular distribution, and spectral characteristics of the resulting $\gamma$-ray emission are consistent with observations, suggesting that disrupted globular clusters could plausibly account for the observed excess in the Galactic Centre.
In the in situ scenario, the Galactic bulge forms as part of the Milky Way disc through internal evolutionary processes (Shen & Li, 2016). In this framework, early star formation is centrally concentrated, and the bulge emerges naturally from the dense, inner regions of the disc. Secular evolution may subsequently lead to the formation of a bar, which can undergo vertical buckling, producing the boxy or peanut-shaped morphology observed today (e.g. Gerhard, 2015; Di Matteo, 2016; Łokas, 2019). This scenario is supported by a range of observational evidence, including a steep radial gradient in stellar ages—with stars in the inner few kiloparsecs typically older than those at larger radii—as well as metallicity and $\alpha$-element abundance patterns that suggest rapid early enrichment (e.g. Zoccali et al., 2014; Bovy et al., 2019; Sit & Ness, 2020). These trends are consistent with a bulge dominated by an old, in situ stellar population that formed early in the Galaxy’s history and evolved alongside the disc.
@@ 3.1 Accreted scenario
Here, we describe a simplified population synthesis model based on the assumption that the Galactic bulge was assembled through the accretion and disruption of multiple globular clusters. In this case, NS formation does not need to occur within the same binary system that later undergoes the LMXB phase. Instead, NSs may acquire new companions through dynamical interactions and subsequently evolve through the LMXB stage. For our purposes, we construct the population synthesis around the Tauris & Savonije (1999) relation, which links orbital period and companion mass for systems that have undergone stable mass transfer. This relation is relatively insensitive to the effects of dynamical evolution within clusters. Notably, MSPs observed in globular clusters follow this relation as well (see Fig. 1).
@@
In this scenario, a NS is born within a globular cluster and remains gravitationally bound to the cluster (see Section 5.1 for implications of natal kicks). At a later time, the NS acquires a low-mass companion (M<2M<2\,M) through dynamical interactions. The resulting binary undergoes stable mass transfer (i.e. an LMXB phase), during which the NS is spun up (i.e. recycled) to become an MSP, and the donor evolves into a He WD. The system is expected to settle onto the Tauris & Savonije (1999) relation at the end of the mass transfer phase. Subsequently, the host globular cluster is accreted and merges with the central Galactic cluster, depositing the binary into the bulge. From this point onward, no further binary formation occurs. This scenario has several advantages. Binaries in which the NS does not undergo stable mass transfer retain long spin periods and exhibit low spin-down luminosities, making them unlikely contributors to the $\gamma$-ray excess observed in the Galactic Centre. The Tauris & Savonije (1999) relation thus characterises the subset of He WDNS binaries that have gone through the LMXB phase and are capable of producing observable MSPs.
@@
Below we present the details of this population synthesis model. Following the estimates of Holst & Hooper (2025), based on the observed $\gamma$-ray luminosity, we consider two benchmark values for the total number of MSPs (isolated and in binaries) in the Galactic bulge: $3.6\times 10^{4}$ and $2\times 10^{5}$. The lower value assumes that bulge MSPs have $\gamma$-ray luminosities similar to those observed in the field and in globular clusters, while the higher value corresponds to the case in which bulge MSPs are intrinsically fainter and therefore must be more numerous in order to produce the same emission. To reflect the possibility that some recycled binaries may be disrupted in the dense environments of globular clusters after the LMXB phase, we assume a binary fraction of $0.5$, consistent with observations of MSP populations in both globular clusters and the field (see Section 2). Thus, we set $N^{\rm bulge}_{\rm MSP}=1.8\times 10^{4}$ and $1\times 10^{5}$ for the number of MSPs in binaries in the bulge.
@@
The initial mass of the NS is drawn from the bimodal normal distribution proposed by Antoniadis et al. (2016), which reflects the observed mass distribution of MSPs. It is described as:
@@
$+\frac{r}{\sqrt{2\pi}\;\sigma_{2}}\exp\left[-\frac{(M_{\mathrm{MSP}}-\mu_{2})^{2}}{2\sigma_{2}^{2}}\right]$ (7)
The constants are as the following $\mu_{1}=1.39$ M and $\mu_{2}=1.81$ M, with standard deviations $\sigma_{1}=0.06$ M and $\sigma_{2}=0.177$ M, and $r=0.4$.
To model the probability density function along the Tauris & Savonije (1999) relation, we explore two approaches. In the first, we draw the companion mass $M_{\mathrm{WD}}$ from a power-law distribution with slope $\gamma=-4$, denoted as $M_{\mathrm{WD}}\sim\Gamma(-4)$. As shown in the lower-left panel of Fig. 1, this choice is consistent with observations. In the second approach, we draw the orbital period from a log-uniform distribution in the range 0.1–100 days, written as $P_{\mathrm{\rm orb}}\sim\log_{10}\mathcal{U}(-1,2)$. Log-uniform distributions are commonly used in binary population synthesis and also appear plausible given current observational constraints (see lower right panel of Fig. 1). In both cases, the model is expected to overproduce MSPs with the lowest WD masses and shortest orbital periods. This is an intended effect, as radio pulsar searches are known to be incomplete in the short-period regime. The sensitivity of radio searches degrades for MSP binaries with orbital periods below 1 day and He WD mass below 0.19 M. Next, we use the Tauris & Savonije (1999) relation to estimate either the orbital period (if we draw $M_{\mathrm{WD}}$) or the WD mass (if we draw $P_{\mathrm{orb}}$).
Next, we randomly pair each of the synthetic WDs with a NS, forming a population of WDNS binaries. This corresponds to the dynamical pairing of a low-mass star with a NS in the dense environment of a globular cluster, prior to the onset of stable mass transfer. None of these synthetic MSP systems would be immediately detectable by LISA, as their post-LMXB orbital periods remain relatively long ($P_{\rm orb}\sim$ hours). However, we expect their orbits to gradually shrink over time due to the gravitational-wave radiation reaction. To model this inspiral, we must assign an age to each system so that we can evolve its orbital parameters from the end of the LMXB phase to the present day. We adopt two approaches for estimating these ages. In the first, we draw ages from a normal distribution centred at 11 Gyr with a standard deviation of 1 Gyr, motivated by Jiaqi et al. (2025). In the second, we use the synthetic catalogue of Chen & Gnedin (2024), randomly assigning to each system the time of accretion of an associated globular cluster. We compute the orbital decay using the analytical prescription of Peters (1964), assuming zero eccentricity. This assumption is justified, as any significant eccentricity is expected to be damped during the LMXB phase due to stable mass transfer (Phinney, 1992). We first convert the orbital period to a semi-major axis using Kepler’s third law:
@@
We then check whether the binary merges within its assigned age by computing the merger time due to gravitational wave emission:
@@
where $\beta$ is given by:
@@
If the binary does not merge within its expected age, we compute the evolved semi-major axis as:
@@
and convert it back to an orbital period for use in our mock catalogue.
Finally, we assume that binaries are spatially distributed according to a spherically symmetric Gaussian density profile, centred on the Galactic Centre. The distribution is given by
@@
where $r$ is the spherical radial distance from the Galactic Centre, and $r_{\rm b}=0.5$ kpc is the characteristic scale radius. This choice provides a simple approximation for the spatial extent of the bulge MSP population. It is sufficient for our purposes, as the detailed spatial structure of the bulge does not significantly affect detectability with LISA (Storck & Church, 2023).
3.2 In situ scenario
In this study, we consider three models from Korol et al. (2024b) that adopt different prescriptions for the common-envelope phase—the phase of rapid orbital shrinkage that occurs when one star expands to fill its Roche lobe and engulfs its companion. Our model ‘in situ $\alpha\alpha$’ adopts the commonly used $\alpha$-formalism, which assumes that a fraction $\alpha$ of the orbital energy is used to eject the envelope, with the parameter $\lambda$ describing the envelope’s binding energy (Webbink, 1984). In this model, we assume a fixed value of $\alpha\lambda=2$. The model ‘in situ $\alpha\alpha 2$’ also uses the $\alpha$-formalism, but with a lower efficiency parameter, $\alpha\lambda=0.25$. In contrast, the model ‘in situ $\gamma\alpha$’ adopts an alternative prescription for the common-envelope phase, known as the $\gamma$-formalism (Nelemans et al., 2000), which is based on the conservation of angular momentum rather than energy. In this model, the $\gamma$-prescription is applied unless the binary contains a compact object or the common-envelope phase is triggered by a tidal instability (rather than dynamically unstable Roche-lobe overflow). As a result, the first common-envelope episode is typically treated using the $\gamma$-formalism, while the second—typically involving a giant donor and a WD or NS companion—is modelled using the $\alpha$-formalism (see Toonen et al., 2012). For this model, we adopt $\alpha\lambda=2$ and $\gamma=1.75$. Our choice of efficiency parameter values is guided by previous studies that calibrated them using observations. The $\alpha\alpha$ and $\gamma\alpha$ models follow the prescriptions constrained by observed double WD systems (hereafter WDWD; e.g., Nelemans et al., 2000; Nelemans & Tout, 2005; van der Sluys et al., 2006), while the lower efficiency in the $\alpha\alpha 2$ model is based on calibration from compact WDs in binaries with M-type main-sequence companions (Zorotovic et al., 2010; Toonen & Nelemans, 2013; Camacho et al., 2014; Zorotovic et al., 2014). In all three models, we adopt the NS natal kick prescription from Verbunt et al. (2017) as a distribution which describes observation of isolated radio pulsars well.
Our Milky Way model follows the approach of Nelemans et al. (2004), with the star formation rate linearly interpolated from the plane-projected grid computed by Boissier & Prantzos (1999), which is based on a spectro-photometric model of Galactic stellar populations. To represent the stellar bulge, we increase the star formation rate by a factor of two within the inner 3 kpc compared to the original Boissier & Prantzos (1999) prescription, and distribute the resulting binary population according to Eq. (12). This yields a present-day integrated stellar mass of approximately $2.6\times 10^{10}\,$M within this region.
4 Results
Figure 3: Detectable MSPWD binaries in the chirp mass–gravitational wave frequency plane for both formation scenarios considered in this study. The in situ population is shown with orange stars, while systems from the accreted scenario are marked with green downward-pointing triangles. For comparison, the underlying population of detectable WDWD binaries is shown in blue-grey. Contours indicate regions of constant relative uncertainty in the chirp mass ($\sigma_{\mathcal{M}}/\mathcal{M}$) at levels of 1, 0.1, and 0.01, highlighting LISA’s ability to constrain binary parameters. The hatched region marks the part of parameter space where the chirp mass cannot be reliably measured ($\sigma_{\mathcal{M}}/\mathcal{M}>1$). The top axis shows the corresponding orbital period.
Table 1: Summary of the population synthesis models and predicted numbers of MSP binaries detectable with LISA. For each model, we list the initial MSP binary and age distributions, the total number of binaries in the bulge ($N_{\mathrm{MSP}}^{\mathrm{bulge}}$), the number of MSPWD systems detectable with LISA ($N_{\mathrm{MSP}}^{\mathrm{LISA}}$), i.e. those with $\rho_{4\mathrm{yr}}>7$, and the detection fraction $f^{\mathrm{LISA}}$, defined as the ratio of detectable systems to the total modelled bulge population. We use $\Gamma(p)$ to denote a power-law with exponent $p$, and $N(\mu,\sigma)$ for a normal distribution.
@@ 4 Results
Accreted C $M_{\mathrm{WD}}\sim\Gamma(-4)$ Chen & Gnedin (2024) 100 000 (18 000) 4 (1) $4\times 10^{-5}$
Accreted D $P_{\mathrm{orb}}\sim\log_{10}\mathrm{U}(-1,2)$ Chen & Gnedin (2024) 100 000 (18 000) 3 (1) $3\times 10^{-5}$

### v2

@@ 2.2 Detection with LISA
While traditional radio pulsar searches are limited in their ability to detect short-period binaries due to orbital acceleration effects (see bottom panel of Fig. 2), LISA will offer a complementary window into the MSP binary population. In particular, LISA will be sensitive to systems with orbital periods below one hour across the entire Milky Way (Amaro-Seoane et al., 2023). In the case of circular binaries, as considered here due to the orbit being circularised during the recycling phase via stable mass transfer, the system emits continuous, quasi-monochromatic gravitational waves. The gravitational-wave signal from such a binary can be described by eight parameters:
@@
where $\mathcal{A}$ is the gravitational-wave amplitude, $f_{\rm GW}=2/P_{\rm orb}$ is the gravitational-wave frequency (twice the orbital frequency), and $\dot{f}_{\rm GW}$ is its time derivative. The parameters $(\lambda,\beta)$ denote the ecliptic longitude and latitude, respectively; $\iota$ is the inclination angle of the orbital plane; $\psi$ is the polarisation angle; and $\phi_{0}$ is the initial phase of the waveform (e.g. LISA Consortium Waveform Working Group et al., 2023). The gravitational-wave amplitude is given by
@@
where $G$ is the gravitational constant, $c$ is the speed of light, and $d$ is the distance to the binary. This shows that the gravitational-wave amplitude depends on the binary’s frequency, distance, and chirp mass ($\mathcal{M}$), defined as
@@
where $T_{\rm obs}$ is the observation time and $S_{n}(f)$ is the LISA noise power spectral density. In the following analysis, we assume the LISA noise power spectral density as given in the LISA Science Requirements Document (Babak et al., 2021). To assess a binary’s detectability, one can invert Eq. (5), setting the observation time and a detection threshold.
Figure 3 illustrates the detectable region in the companion mass–-gravitational-wave frequency parameter space, assuming an MSP mass of $1.4\penalty 10000\ \mathrm{M}_{\odot}$ and distances of 5 (dark grey), 8 (grey), and 11 kpc (light grey). These distances span the inner Milky Way, which is relevant for the observed $\gamma$-ray excess (e.g. Hooper and Slatyer, 2013). The figure shows that LISA detectability at these distances requires frequencies of at least $\sim$1 mHz (corresponding to orbital periods of less than $\sim$30 min). For lower companion masses, the required frequency increases: approximately 3 mHz for $\sim$0.1 M, and up to $\sim$$10^{-2}$ Hz for $\sim$0.02 M.
Figure 3: Detectability of a binary system containing a 1.4 M MSP with LISA, shown as a function of companion mass and gravitational-wave frequency (bottom x-axis) or orbital period (top x-axis). Grey contours indicate LISA’s 4-year signal-to-noise threshold ($\rho_{\rm 4\,yr}=7$) for sources located at 5 (dark grey), 8 (grey), and 11 kpc (light grey). Dashed horizontal lines mark approximate boundaries between different classes of MSP binaries: black widows (low-mass companions, $M\lesssim 0.06$ M), redbacks and He WDs (intermediate-mass companions, $0.06\lesssim M\lesssim 0.5$ M), and more massive donors such as CO/ONeMg WDs or He stars ($M\gtrsim 0.5$ M). The grey solid curve marks the Roche-lobe overflow (RLOF) boundary for WD donors, indicating the onset of stable mass transfer. Observed MSP binaries are shown as coloured symbols, following the same classification as in Fig. 1. UCXBs lying near the WD RLOF line are shown as light blue pentagons. The two-segment grey arrow illustrates the expected evolutionary track of close binaries.
While the detection of NS binaries in the inner 3 kpc is, in principle, possible at gravitational-wave frequencies above a few millihertz, the companion star may—depending on its mass—fill its Roche lobe and initiate mass transfer before the system reaches the higher-frequency part of the LISA sensitivity window. As a reference, we plot in Fig. 3 the Roche lobe–filling radius of a WD donor using the commonly adopted zero-temperature radius–mass relation from Verbunt and Rappaport (1988):
@@
Binary evolution simulations show that MSPs with He WD companions initiate mass transfer at orbital periods of 5–7 minutes (e.g. Sengar et al., 2017), after which the orbit begins to widen as the donor expands in response to mass loss. This sets an upper frequency limit for detecting such systems in the detached phase. In contrast, simulations by Bobrick et al. (2017) show that systems with CO WD or ONeMg WD companions undergo unstable mass transfer when reaching these short orbital periods, leading to the tidal disruption of the donor. Systems with He-star companions—i.e. stripped stellar cores still burning helium and thus larger and less compact than degenerate WDs—with masses up to approximately $1.2\,\mathrm{M}_{\odot}$ can initiate stable mass transfer at orbital periods between 10 and 80 minutes, essentially set by the Roche lobe radius of the He star (Götberg et al., 2020; Wang et al., 2021). The exact minimum period depends on the He star’s mass, with shorter periods corresponding to physically smaller (lower-mass) donors. Similar to systems with He WD donors, these binaries evolve towards longer orbital periods after reaching the minimum. This characteristic evolution—initial inspiral driven by gravitational-wave emission, followed by mass transfer and orbital widening—is schematically illustrated with a grey two-segment arrow in Fig. 3. More massive He stars (above $1.2\,\mathrm{M}_{\odot}$) tend to initiate mass transfer that is too rapid to remain stable, often resulting in unstable mass transfer or mergers.
@@ 3 Population synthesis of MSP binaries in the Milky Way bulge
Many theories have been proposed to explain the origin of the Milky Way’s bulge (e.g. Nataf, 2017; Shen and Li, 2016), each with distinct implications for the underlying stellar populations and binary formation channels. Here, we consider two contrasting scenarios: one in which the bulge formed in situ as part of the Milky Way disc through internal evolutionary processes, and another in which it assembled through the accretion and disruption of globular clusters. While these scenarios may be somewhat simplified, they provide a useful framework for exploring the potential properties and detectability of MSP binary populations in the bulge.
In the accretion scenario, the main idea is that, due to dynamical friction, globular clusters gradually spiral towards the Galactic centre. In the inner few kiloparsecs, the tidal field of the Galaxy can exceed a cluster’s self-gravity, stripping stars and potentially dissolving the cluster entirely (e.g. Tremaine et al., 1975; Capuzzo-Dolcetta, 1993; Antonini et al., 2012; Arca-Sedda and Capuzzo-Dolcetta, 2014; Gnedin et al., 2014). The globular clusters observed today are therefore likely the survivors of this process, representing only a small fraction of the original population. Globular clusters in the Milky Way halo are also known to be overabundant in MSPs (see Section 2). It is therefore reasonable to hypothesise that the MSPs responsible for the $\gamma$-ray excess could have originated in dense star clusters in the halo, which were subsequently disrupted, depositing their MSPs into the Galactic bulge. Brandt and Kocsis (2015) showed that, in this scenario, the amplitude, angular distribution, and spectral characteristics of the resulting $\gamma$-ray emission are consistent with observations, suggesting that disrupted globular clusters could plausibly account for the observed excess in the Galactic centre.
In the in situ scenario, the Galactic bulge forms as part of the Milky Way disc through internal evolutionary processes (Shen and Li, 2016). In this framework, early star formation is centrally concentrated, and the bulge emerges naturally from the dense, inner regions of the disc. Secular evolution may subsequently lead to the formation of a bar, which can undergo vertical buckling, producing the boxy or peanut-shaped morphology observed today (e.g. Gerhard, 2015; Di Matteo, 2016; Łokas, 2019). This scenario is supported by a range of observational evidence, including a steep radial gradient in stellar ages—with stars in the inner few kiloparsecs typically older than those at larger radii—as well as metallicity and $\alpha$-element abundance patterns that suggest rapid early enrichment (e.g. Zoccali et al., 2014; Bovy et al., 2019; Sit and Ness, 2020). These trends are consistent with a bulge dominated by an old, in situ stellar population that formed early in the Galaxy’s history and evolved alongside the disc.
@@ 3.1 Accreted scenario
Here, we describe a simplified population synthesis model based on the assumption that the Galactic bulge was assembled through the accretion and disruption of multiple globular clusters. In this case, NS formation does not need to have occurred within the same binary system that later undergoes the LMXB phase. Instead, NSs may acquire new companions through dynamical interactions and subsequently evolve through the LMXB stage. For our purposes, we construct the population synthesis around the Tauris and Savonije (1999) relation, which links orbital period and companion mass for systems that have undergone stable mass transfer. This relation is relatively insensitive to the effects of dynamical evolution within clusters. Notably, MSPs observed in globular clusters follow this relation as well (see Fig. 1).
@@
In this scenario, a NS is born within a globular cluster and remains gravitationally bound to the cluster (see Section 5.2 for implications of natal kicks). At a later time, the NS acquires a low-mass companion (M<2M<2\,M) through dynamical interactions. The resulting binary undergoes stable mass transfer (i.e. an LMXB phase), during which the NS is spun up (i.e. recycled) to become an MSP, and the donor evolves into a He WD. The system is expected to settle onto the Tauris and Savonije (1999) relation at the end of the mass transfer phase. Subsequently, the host globular cluster is accreted and merges with the central Galactic cluster, depositing the binary into the bulge. From this point onward, no further binary formation occurs. This scenario has several advantages. Binaries in which the NS does not undergo stable mass transfer retain long spin periods and exhibit low spin-down luminosities, making them unlikely contributors to the $\gamma$-ray excess observed in the Galactic centre. The Tauris and Savonije (1999) relation thus characterises the subset of He WDNS binaries that have gone through the LMXB phase and are capable of producing observable MSPs.
@@
Below we present the details of this population synthesis model. Following the estimates of Holst and Hooper (2025), based on the observed $\gamma$-ray luminosity, we consider two benchmark values for the total number of MSPs (isolated and in binaries) in the Galactic bulge: $3.6\times 10^{4}$ and $2\times 10^{5}$. The lower value assumes that bulge MSPs have $\gamma$-ray luminosities similar to those observed in the field and in globular clusters, while the higher value corresponds to the case in which bulge MSPs are intrinsically fainter and therefore must be more numerous in order to produce the same emission. To reflect the possibility that some recycled binaries may be disrupted in the dense environments of globular clusters after the LMXB phase, we assume a binary fraction of $0.5$, consistent with observations of MSP populations in both globular clusters and the field (see Section 2). Thus, we set $N^{\rm bulge}_{\rm MSP}=1.8\times 10^{4}$ and $1\times 10^{5}$ for the number of MSPs in binaries in the bulge.
@@
The initial mass of the NS is drawn from the bimodal normal distribution proposed by Antoniadis et al. (2016), which reflects the observed mass distribution of MSPs. It is described as
@@
$+\frac{r}{\sqrt{2\pi}\;\sigma_{2}}\exp\left[-\frac{(M_{\mathrm{MSP}}-\mu_{2})^{2}}{2\sigma_{2}^{2}}\right].$ (7)
The constants are as follows: $\mu_{1}=1.39$ M and $\mu_{2}=1.81$ M, with standard deviations $\sigma_{1}=0.06$ M and $\sigma_{2}=0.177$ M, and $r=0.4$.
To model the probability density function along the Tauris and Savonije (1999) relation, we explored two approaches. In the first, we drew the companion mass $M_{\mathrm{WD}}$ from a power-law distribution with slope $\gamma=-4$, denoted as $M_{\mathrm{WD}}\sim\Gamma(-4)$. As shown in the top panel of Fig. 2, this choice is consistent with observations. In the second approach, we drew the orbital period from a log-uniform distribution in the range 0.1–100 days, written as $P_{\mathrm{\rm orb}}\sim\log_{10}\mathcal{U}(-1,2)$. Log-uniform distributions are commonly used in binary population synthesis and also appear plausible given current observational constraints (see bottom panel of Fig. 2). In both cases, the model is expected to overproduce MSPs with the lowest WD masses and shortest orbital periods. This is an intended effect, as radio pulsar searches are known to be incomplete in the short-period regime. The sensitivity of radio searches degrades for MSP binaries with orbital periods below 1 day and He WD mass below 0.19 M. Next, we use the Tauris and Savonije (1999) relation to estimate either the orbital period (if we draw $M_{\mathrm{WD}}$) or the WD mass (if we draw $P_{\mathrm{orb}}$).
Next, we randomly paired each of the synthetic WDs with a NS, forming a population of WDNS binaries. This corresponds to the dynamical pairing of a low-mass star with a NS in the dense environment of a globular cluster, prior to the onset of stable mass transfer. None of these synthetic MSP systems would be immediately detectable by LISA, as their post-LMXB orbital periods remain relatively long ($P_{\rm orb}$ of approximately a few hours). However, we expect their orbits to gradually shrink over time due to the gravitational-wave radiation reaction. To model this inspiral, we must assign an age to each system so that we can evolve its orbital parameters from the end of the LMXB phase to the present day. We adopt two approaches for estimating these ages. In the first, we draw ages from a normal distribution centred at 11 Gyr with a standard deviation of 1 Gyr, motivated by Ying et al. (2025). In the second, we use the synthetic catalogue of Chen and Gnedin (2024), randomly assigning to each system the time of accretion of an associated globular cluster. We computed the orbital decay using the analytical prescription of Peters (1964), assuming zero eccentricity. This assumption is justified, as any significant eccentricity is expected to be damped during the LMXB phase due to stable mass transfer (Phinney, 1992). We first converted the orbital period to a semi-major axis using Kepler’s third law:
@@
We then checked whether the binary merges within its assigned age by computing the merger time due to gravitational-wave emission:
@@
where $\beta$ is given by
@@
If the binary does not merge within its expected age, we computed the evolved semi-major axis as
@@
and converted it back to an orbital period for use in our mock catalogue.
Finally, we assumed that binaries are spatially distributed according to a spherically symmetric Gaussian density profile, centred on the Galactic centre. The distribution is given by
@@
where $r$ is the spherical radial distance from the Galactic centre, and $r_{\rm b}=0.5$ kpc is the characteristic scale radius. This choice provides a simple approximation for the spatial extent of the bulge MSP population. It is sufficient for our purposes, as the detailed spatial structure of the bulge does not significantly affect detectability with LISA (Storck and Church, 2023).
Table 1: Summary of the population synthesis models and predicted numbers of MSP binaries detectable with LISA.  [^3] For each model, we list the initial MSP binary and age distributions, the total number of binaries in the bulge ($N_{\mathrm{MSP}}^{\mathrm{bulge}}$), the number of MSPWD systems detectable with LISA ($N_{\mathrm{MSP}}^{\mathrm{LISA}}$), i.e. those with $\rho_{4\mathrm{yr}}>7$, and the detection fraction $f^{\mathrm{LISA}}$, defined as the ratio of detectable systems to the total modelled bulge population. We use $\Gamma(p)$ to denote a power-law with exponent $p$, and $N(\mu,\sigma)$ for a normal distribution
@@
Accreted C $M_{\mathrm{WD}}\sim\Gamma(-4)$ Chen and Gnedin (2024) 100 000 (18 000) 4 (1) $4\times 10^{-5}$
Accreted D $P_{\mathrm{orb}}\sim\log_{10}\mathrm{U}(-1,2)$ Chen and Gnedin (2024) 100 000 (18 000) 3 (1) $3\times 10^{-5}$